Clark Kent goes to the bakery every morning. He jogs 300m north and 400m west of his house. After eating his morning donut ,he changes into superman and flies home. If he flies straight home using the shortest path possible, how far does superman have to fly get home??????????? how?
Answer by rfer(16322)

c^2=300^2+400^2
c^2=90000+160000
c^2=250000
c=sqrt 250000
c=500 m